### v2

@@ A.1.2 Some technical lemmas
Under the assumptions in Theorem 3.1, the solution $u(\mathbf{x})$ to the Poisson equation Eq. 31 exists, unique up to a constant, and satisfies the gradient estimate
@@
$\mathcal{L}u(\mathbf{x}):=\Delta u(\mathbf{x})+\nabla\log\pi_{\varepsilon}(\mathbf{x})\cdot\nabla u(\mathbf{x}).$
@@
For simplicity, we still denote $\mathcal{L}\mathbf{v}(\mathbf{x})$ when $\mathbf{v}$ is a vector-valued function
$\mathcal{L}\mathbf{v}(\mathbf{x}):=(\mathcal{L}\mathbf{v}_{1}(\mathbf{x}),\dots\mathcal{L}\mathbf{v}_{b}(\mathbf{x}))^{\rm T}.$
@@
It suffices to prove Eq. 34 for $\phi\in C^{1}\cap{\rm Lip}_{1}^{0}(\pi_{\varepsilon,i})$. Note this space is dense in ${\rm Lip}_{1}^{0}(\pi_{\varepsilon,i})$, so for general $\phi\in{\rm Lip}_{1}(\pi_{\varepsilon,i})$, we can take a sequence of $\phi_{k}\in C^{1}\cap{\rm Lip}_{1}^{0}(\pi_{\varepsilon,i})$ that converges to $\phi$. Eq. 34 holds uniformly for $\phi_{k}$, so that passing the limit shows that it holds for any $\phi\in{\rm Lip}_{1}^{0}(\pi_{\varepsilon,i})$. Now fix any $\phi\in C^{1}\cap{\rm Lip}_{1}^{0}(\pi_{\varepsilon,i})$. It is straightforward to verify by standard elliptic theory that the solution exists (up to a constant) and $u\in C^{3}$. Taking gradient w.r.t. $\mathbf{x}_{j}$ in Eq. 31, we obtain
(35) $\nabla_{j}\phi(\mathbf{x}_{i})=\mathcal{L}(\nabla_{j}u)(\mathbf{x})+\sum_{k}\nabla_{jk}^{2}\log\pi_{\varepsilon}(\mathbf{x})\nabla_{k}u(\mathbf{x}).$
@@
Recall by definition Eq. 14, $\pi_{\varepsilon}(\mathbf{x})\propto\exp(-l(\mathbf{x})-\varphi_{\varepsilon}(\mathbf{x}))$. Direct computation shows
$\nabla^{2}\log\pi_{\varepsilon}(\mathbf{x})=-\nabla^{2}l(\mathbf{x})-\nabla^{2}\varphi_{\varepsilon}(\mathbf{x})=-\lambda\mathbf{A}^{\rm T}\mathbf{A}-\nabla^{2}\varphi_{\varepsilon}(\mathbf{x}),$
where the $(i,j)$-th subblock of $\nabla^{2}\varphi_{\varepsilon}(\mathbf{x})$ is given by
@@
(36) $\begin{split}&\nabla_{ij}^{2}\varphi_{\varepsilon}(\mathbf{x})=\delta\sum_{\alpha=1}^{d}\left((\mathbf{D}_{\alpha}^{(v)}\mathbf{x})^{2}+(\mathbf{D}_{\alpha}^{(h)}\mathbf{x})^{2}+\varepsilon\right)^{-3/2}\\ &\cdot\Big{[}\left((\mathbf{D}_{\alpha}^{(v)}\mathbf{x})\mathbf{D}_{\alpha,i}^{(v)}-(\mathbf{D}_{\alpha}^{(h)}\mathbf{x})\mathbf{D}_{\alpha,i}^{(h)}\right)^{\rm T}\left((\mathbf{D}_{\alpha}^{(v)}\mathbf{x})\mathbf{D}_{\alpha,j}^{(v)}-(\mathbf{D}_{\alpha}^{(h)}\mathbf{x})\mathbf{D}_{\alpha,j}^{(h)}\right)\\ &\quad+\varepsilon\left((\mathbf{D}_{\alpha,i}^{(v)})^{\rm T}\mathbf{D}_{\alpha,j}^{(v)}+(\mathbf{D}_{\alpha,i}^{(h)})^{\rm T}\mathbf{D}_{\alpha,j}^{(h)}\right)\Big{]}.\end{split}$
@@
For simplicity, denote $H(\mathbf{x})=\nabla^{2}\varphi_{\varepsilon}(\mathbf{x})$. Multiplying Eq. 35 from left by $-\left(\nabla_{j}u(\mathbf{x})\right)^{\rm T}$,
(37) $\begin{split}-\left(\nabla_{j}u(\mathbf{x})\right)^{\rm T}\nabla_{j}\phi(\mathbf{x}_{i})+&\left(\nabla_{j}u(\mathbf{x})\right)^{\rm T}\mathcal{L}(\nabla_{j}u)(\mathbf{x})\\ =~{}&\sum_{k}\left(\nabla_{j}u(\mathbf{x})\right)^{\rm T}(\lambda\mathbf{A}^{\rm T}\mathbf{A}+H(\mathbf{x}))_{jk}\nabla_{k}u(\mathbf{x}).\end{split}$
@@
Since $\mathbf{A}^{\rm T}\mathbf{A}$ is $c$-diagonal block dominant with matrix $M\in\mathbb{R}^{b\times b}$, and $H_{jj}$ is positive definite since $\varphi_{\varepsilon}$ is convex, so that,
$\sum_{k}\left(\nabla_{j}u(\mathbf{x})\right)^{\rm T}(\lambda\mathbf{A}^{\rm T}\mathbf{A}+H(\mathbf{x}))_{jk}\nabla_{k}u(\mathbf{x})$
@@
Next we control $\|H_{jk}(\mathbf{x})\|_{2}$. Note $\mathbf{D}^{(v)}_{\alpha,i}=0$ if $\alpha$ does not live in block $i$ or a vertical neighbor of block $i$. Similarly for $\mathbf{D}^{(h)}_{\alpha,i}$. Therefore, $H_{ij}(\mathbf{x})=0$ if $i$ and $j$ are not neighbors. For vertical neighbors $i,j$, there are exactly $2m$ boundary indices $\alpha$ s.t. $\left(\mathbf{D}^{(v)}_{\alpha,i}\right)^{\rm T}\mathbf{D}^{(v)}_{\alpha,j}\neq 0$, and $\left\|\left(\mathbf{D}^{(v)}_{\alpha,i}\right)^{\rm T}\mathbf{D}^{(v)}_{\alpha,j}\right\|_{2}\leq 2$. Moreover,
$\left\|\left((\mathbf{D}_{\alpha}^{(v)}\mathbf{x})\mathbf{D}_{\alpha,i}^{(v)}-(\mathbf{D}_{\alpha}^{(h)}\mathbf{x})\mathbf{D}_{\alpha,i}^{(h)}\right)^{\rm T}\left((\mathbf{D}_{\alpha}^{(v)}\mathbf{x})\mathbf{D}_{\alpha,j}^{(v)}-(\mathbf{D}_{\alpha}^{(h)}\mathbf{x})\mathbf{D}_{\alpha,j}^{(h)}\right)\right\|_{2}$
$\leq\left(\sqrt{2}|\mathbf{D}_{\alpha}^{(v)}\mathbf{x}|+\sqrt{2}|\mathbf{D}_{\alpha}^{(h)}\mathbf{x}|\right)^{2}\leq 4\left((\mathbf{D}_{\alpha}^{(v)}\mathbf{x})^{2}+(\mathbf{D}_{\alpha}^{(h)}\mathbf{x})^{2}\right)$
Therefore, recall Eq. 36, it holds that
(38) $\begin{split}\left\|H_{ij}(\mathbf{x})\right\|_{2}\leq~{}&\delta\sum_{\alpha\text{ in boundary}}\left((\mathbf{D}_{\alpha}^{(v)}\mathbf{x})^{2}+(\mathbf{D}_{\alpha}^{(h)}\mathbf{x})^{2}+\varepsilon\right)^{-3/2}\\ &\cdot\left[4\left((\mathbf{D}_{\alpha}^{(v)}\mathbf{x})^{2}+(\mathbf{D}_{\alpha}^{(h)}\mathbf{x})^{2}\right)+4\varepsilon\right]\\ \leq~{}&4\delta\sum_{\alpha\text{ in boundary}}\left((\mathbf{D}_{\alpha}^{(v)}\mathbf{x})^{2}+(\mathbf{D}_{\alpha}^{(h)}\mathbf{x})^{2}+\varepsilon\right)^{-1/2}\leq 8m\delta\varepsilon^{-1/2}.\end{split}$
Similar result holds when $i,j$ are horizontal neighbors. Also notice $H_{jk}(\mathbf{x})=0$ when $j\not\sim k$, where we denote $j\sim k$ if $j\neq k$ are neighboring blocks. Now substitute the above inequalities into Eq. 37, we have
(39) $\begin{split}&-\left(\nabla_{j}u(\mathbf{x})\right)^{\rm T}\nabla_{j}\phi(\mathbf{x}_{i})+\left(\nabla_{j}u(\mathbf{x})\right)^{\rm T}\mathcal{L}(\nabla_{j}u)(\mathbf{x})\\ \geq~{}&\lambda M_{jj}\|\nabla_{j}u(\mathbf{x})\|_{2}^{2}-\sum_{k\neq j}(\lambda M_{jk}+8m\delta\varepsilon^{-1/2}{\bf{1}}_{j\sim k})\|\nabla_{j}u(\mathbf{x})\|_{2}\|\nabla_{k}u(\mathbf{x})\|_{2}.\end{split}$
Consider $\mathbf{x}$ where $\|\nabla_{j}u(\mathbf{x})\|_{2}$ reaches its maximum, i.e. $\|\nabla_{j}u(\mathbf{x})\|_{2}=\left\|\nabla_{j}u\right\|_{L^{\infty}}$. The first order optimality condition reads
$0=\nabla\left(\|\nabla_{j}u(\mathbf{x})\|_{2}^{2}\right)=\nabla\nabla_{j}u(\mathbf{x})\cdot\nabla_{j}u(\mathbf{x}),$
@@
$0\geq\Delta\left(\|\nabla_{j}u(\mathbf{x})\|_{2}^{2}\right)=2\|\nabla\nabla_{j}u(\mathbf{x})\|_{\rm F}^{2}+2(\nabla_{j}u(\mathbf{x}))^{\rm T}\Delta\nabla_{j}u(\mathbf{x}).$
So that $(\nabla_{j}u(\mathbf{x}))^{\rm T}\Delta\nabla_{j}u(\mathbf{x})\leq 0$ since $\|\nabla\nabla_{j}u(\mathbf{x})\|_{\rm F}^{2}\geq 0$. Under these conditions,
$(\nabla_{j}u(\mathbf{x}))^{\rm T}\mathcal{L}(\nabla_{j}u(\mathbf{x}))=(\nabla_{j}u(\mathbf{x}))^{\rm T}\Delta\nabla_{j}u(\mathbf{x})+\nabla\log\pi_{\varepsilon}(\mathbf{x})\cdot\nabla\nabla_{j}u(\mathbf{x})\cdot\nabla_{j}u(\mathbf{x})\leq 0.$
So that at the maximum point, Eq. 39 reads
$\lambda M_{jj}\|\nabla_{j}u\|_{L^{\infty}}^{2}-$ $\sum_{k\neq j}(\lambda M_{jk}+8m\delta\varepsilon^{-1/2}{\bf{1}}_{j\sim k})\left\|\nabla_{j}u\right\|_{L^{\infty}}\|\nabla_{k}u(\mathbf{x})\|_{2}$
@@
Here we use $\|\nabla_{j}\phi(\mathbf{x}_{i})\|_{2}\leq\delta_{ij}$, since $\phi$ is $1$-Lipschitz and is only a function of $\mathbf{x}_{i}$. So if $\|\nabla_{j}u\|_{L^{\infty}}>0$, it holds that
@@
$\max_{\alpha}\mathbb{E}_{\mathbf{x}\sim\pi}\left\|\mathbf{D}_{\alpha}\mathbf{x}\right\|_{2}^{-1}\leq C_{\pi}.$
@@
$\mathbf{D}_{\alpha}\mathbf{x}=(\mathbf{D}_{\alpha}^{(v)}\mathbf{x},\mathbf{D}_{\alpha}^{(h)}\mathbf{x})=(\mathbf{x}_{\alpha}-\mathbf{x}_{\alpha}^{(v)},\mathbf{x}_{\alpha}-\mathbf{x}_{\alpha}^{(h)}).$
Introduce change of variable $\mathbf{z}=T\mathbf{x}$ for some linear map determined via
$\mathbf{z}_{\alpha}^{(v)}=\mathbf{x}_{\alpha}^{(v)}-\mathbf{x}_{\alpha},\quad\mathbf{z}_{\alpha}^{(h)}=\mathbf{x}_{\alpha}^{(h)}-\mathbf{x}_{\alpha},$
@@
and $\mathbf{z}_{\alpha-}=\mathbf{x}_{\alpha-}\in\mathbb{R}^{d-2}$ for the other coordinates. Accordingly $\pi(\mathbf{x})$ is transformed into another distribution $\mu(\mathbf{z})=\pi(T^{-1}\mathbf{z})$ (note $\det T=1$). Also note that $T^{-1}$ admits explicit form, i.e.
$\mathbf{x}_{\alpha}^{(v)}=\mathbf{z}_{\alpha}^{(v)}+\mathbf{z}_{\alpha},\quad\mathbf{x}_{\alpha}^{(h)}=\mathbf{z}_{\alpha}^{(h)}+\mathbf{z}_{\alpha},\quad\mathbf{x}_{\alpha-}=\mathbf{z}_{\alpha-}.$
Denote $\mathbf{z}_{\alpha+}=(\mathbf{z}_{\alpha}^{(v)},\mathbf{z}_{\alpha}^{(h)})$ for convenience. Consider the factorization
$\mu(\mathbf{z})=\mu(\mathbf{z}_{\alpha+},\mathbf{z}_{\alpha-})=\mu(\mathbf{z}_{\alpha+}|\mathbf{z}_{\alpha-})\mu(\mathbf{z}_{\alpha-}).$
where $\mu(\mathbf{z}_{\alpha-})$ denotes the marginal of $\mu$ on $\mathbf{z}_{\alpha-}$. Notice
@@
$\log\mu(\mathbf{z}_{\alpha+}|\mathbf{z}_{\alpha-})=-\frac{\lambda}{2}\left\|\mathbf{y}-\mathbf{A}T^{-1}\mathbf{z}\right\|_{2}^{2}-\delta\sum_{\alpha=1}^{d}\left\|\mathbf{D}_{\alpha}(T^{-1}\mathbf{z})\right\|_{2}-\log\mu(\mathbf{z}_{\alpha-})+\text{const}.$
@@
Fix $\mathbf{z}_{\alpha-}$ for the moment. Notice $\frac{\lambda}{2}\left\|\mathbf{y}-\mathbf{A}T^{-1}\mathbf{z}\right\|_{2}^{2}$ is $L$-smooth for some dimension-independent $L>0$, since only a dimension-independent number of coordinates of $\mathbf{A}T^{-1}\mathbf{z}$ depend on $\mathbf{z}_{\alpha-}$. Therefore, fix any $\mathbf{z}_{\alpha+}^{0}$,
$\frac{\lambda}{2}\left\|\mathbf{y}-\mathbf{A}T^{-1}(\mathbf{z}_{\alpha+},\mathbf{z}_{\alpha-})\right\|_{2}^{2}-\frac{\lambda}{2}\left\|\mathbf{y}-\mathbf{A}T^{-1}(\mathbf{z}_{\alpha+}^{0},\mathbf{z}_{\alpha-})\right\|_{2}^{2}-\mathbf{v}^{0}\cdot(\mathbf{z}_{\alpha+}-\mathbf{z}_{\alpha+}^{0})$
$\leq\frac{L}{2}\left\|\mathbf{z}_{\alpha+}-\mathbf{z}_{\alpha+}^{0}\right\|_{2}^{2}$ ,
@@
where $\mathbf{v}^{0}$ is the gradient w.r.t. $\mathbf{z}_{\alpha+}$ of $\frac{\lambda}{2}\left\|\mathbf{y}-\mathbf{A}T^{-1}\mathbf{z}\right\|_{2}^{2}$ at $\mathbf{z}_{\alpha+}^{0}$. Notice also
$\delta\sum_{\alpha=1}^{d}\left\|\mathbf{D}_{\alpha}(T^{-1}(\mathbf{z}_{\alpha+},\mathbf{z}_{\alpha-}))\right\|_{2}\leq\delta\sum_{\alpha=1}^{d}\left\|\mathbf{D}_{\alpha}(T^{-1}(\mathbf{z}_{\alpha+}^{0},\mathbf{z}_{\alpha-}))\right\|_{2}+8\delta\left\|\mathbf{z}_{\alpha+}-\mathbf{z}_{\alpha+}^{0}\right\|_{2},$

### v3

@@ A.1.2 Some technical lemmas
Under the assumptions in Theorem 3.1, the solution $u({\bm{x}})$ to the Poisson equation Eq. 31 exists, unique up to a constant, and satisfies the gradient estimate
@@
$\mathcal{L}u({\bm{x}}):=\Delta u({\bm{x}})+\nabla\log\pi_{\varepsilon}({\bm{x}})\cdot\nabla u({\bm{x}}).$
@@
For simplicity, we still denote $\mathcal{L}{\bm{v}}({\bm{x}})$ when ${\bm{v}}$ is a vector-valued function
$\mathcal{L}{\bm{v}}({\bm{x}}):=(\mathcal{L}{\bm{v}}_{1}({\bm{x}}),\dots\mathcal{L}{\bm{v}}_{b}({\bm{x}}))^{\rm T}.$
@@
It suffices to prove Eq. 34 for $\phi\in C^{1}\cap{\rm Lip}_{1}^{0}(\pi_{\varepsilon,i})$. Note this space is dense in ${\rm Lip}_{1}^{0}(\pi_{\varepsilon,i})$, so for general $\phi\in{\rm Lip}_{1}(\pi_{\varepsilon,i})$, we can take a sequence of $\phi_{k}\in C^{1}\cap{\rm Lip}_{1}^{0}(\pi_{\varepsilon,i})$ that converges to $\phi$. Eq. 34 holds uniformly for $\phi_{k}$, so that passing the limit shows that it holds for any $\phi\in{\rm Lip}_{1}^{0}(\pi_{\varepsilon,i})$. Now fix any $\phi\in C^{1}\cap{\rm Lip}_{1}^{0}(\pi_{\varepsilon,i})$. It is straightforward to verify by standard elliptic theory that the solution exists (up to a constant) and $u\in C^{3}$. Taking the gradient w.r.t. ${\bm{x}}_{j}$ in Eq. 31, we obtain
(35) $\nabla_{j}\phi({\bm{x}}_{i})=\mathcal{L}(\nabla_{j}u)({\bm{x}})+\sum_{k}\nabla_{jk}^{2}\log\pi_{\varepsilon}({\bm{x}})\nabla_{k}u({\bm{x}}).$
@@
Recall by definition Eq. 14, $\pi_{\varepsilon}({\bm{x}})\propto\exp(-l({\bm{x}})-\varphi_{\varepsilon}({\bm{x}}))$. Direct computation shows
$\nabla^{2}\log\pi_{\varepsilon}({\bm{x}})=-\nabla^{2}l({\bm{x}})-\nabla^{2}\varphi_{\varepsilon}({\bm{x}})=-\lambda{\bm{A}}^{\rm T}{\bm{A}}-\nabla^{2}\varphi_{\varepsilon}({\bm{x}}),$
where the $(i,j)$-th subblock of $\nabla^{2}\varphi_{\varepsilon}({\bm{x}})$ is given by
@@
(36) $\begin{split}&\nabla_{ij}^{2}\varphi_{\varepsilon}({\bm{x}})=\delta\sum_{\alpha=1}^{d}\left(({\bm{D}}_{\alpha}^{(v)}{\bm{x}})^{2}+({\bm{D}}_{\alpha}^{(h)}{\bm{x}})^{2}+\varepsilon\right)^{-3/2}\\ &\cdot\Big{[}\left(({\bm{D}}_{\alpha}^{(v)}{\bm{x}}){\bm{D}}_{\alpha,i}^{(v)}-({\bm{D}}_{\alpha}^{(h)}{\bm{x}}){\bm{D}}_{\alpha,i}^{(h)}\right)^{\rm T}\left(({\bm{D}}_{\alpha}^{(v)}{\bm{x}}){\bm{D}}_{\alpha,j}^{(v)}-({\bm{D}}_{\alpha}^{(h)}{\bm{x}}){\bm{D}}_{\alpha,j}^{(h)}\right)\\ &\quad+\varepsilon\left(({\bm{D}}_{\alpha,i}^{(v)})^{\rm T}{\bm{D}}_{\alpha,j}^{(v)}+({\bm{D}}_{\alpha,i}^{(h)})^{\rm T}{\bm{D}}_{\alpha,j}^{(h)}\right)\Big{]}.\end{split}$
@@
For simplicity, denote $H({\bm{x}})=\nabla^{2}\varphi_{\varepsilon}({\bm{x}})$. Multiplying Eq. 35 from left by $-\left(\nabla_{j}u({\bm{x}})\right)^{\rm T}$,
(37) $\begin{split}-\left(\nabla_{j}u({\bm{x}})\right)^{\rm T}\nabla_{j}\phi({\bm{x}}_{i})+&\left(\nabla_{j}u({\bm{x}})\right)^{\rm T}\mathcal{L}(\nabla_{j}u)({\bm{x}})\\ =~{}&\sum_{k}\left(\nabla_{j}u({\bm{x}})\right)^{\rm T}(\lambda{\bm{A}}^{\rm T}{\bm{A}}+H({\bm{x}}))_{jk}\nabla_{k}u({\bm{x}}).\end{split}$
@@
Since ${\bm{A}}^{\rm T}{\bm{A}}$ is $c$-diagonal block dominant with matrix $M\in\mathbb{R}^{b\times b}$, and $H_{jj}$ is positive definite, since $\varphi_{\varepsilon}$ is convex, we have
$\sum_{k}\left(\nabla_{j}u({\bm{x}})\right)^{\rm T}(\lambda{\bm{A}}^{\rm T}{\bm{A}}+H({\bm{x}}))_{jk}\nabla_{k}u({\bm{x}})$
@@
Next we control $\|H_{jk}({\bm{x}})\|_{2}$. Note ${\bm{D}}^{(v)}_{\alpha,i}=0$ if $\alpha$ does not live in block $i$ or a vertical neighbor of block $i$. Similarly for ${\bm{D}}^{(h)}_{\alpha,i}$. Therefore, $H_{ij}({\bm{x}})=0$ if $i$ and $j$ are not neighbors. For vertical neighbors $i,j$, there are exactly $2m$ boundary indices $\alpha$ s.t. $\left({\bm{D}}^{(v)}_{\alpha,i}\right)^{\rm T}{\bm{D}}^{(v)}_{\alpha,j}\neq 0$, and $\left\|\left({\bm{D}}^{(v)}_{\alpha,i}\right)^{\rm T}{\bm{D}}^{(v)}_{\alpha,j}\right\|_{2}\leq 2$. Moreover,
$\left\|\left(({\bm{D}}_{\alpha}^{(v)}{\bm{x}}){\bm{D}}_{\alpha,i}^{(v)}-({\bm{D}}_{\alpha}^{(h)}{\bm{x}}){\bm{D}}_{\alpha,i}^{(h)}\right)^{\rm T}\left(({\bm{D}}_{\alpha}^{(v)}{\bm{x}}){\bm{D}}_{\alpha,j}^{(v)}-({\bm{D}}_{\alpha}^{(h)}{\bm{x}}){\bm{D}}_{\alpha,j}^{(h)}\right)\right\|_{2}$
$\leq\left(\sqrt{2}|{\bm{D}}_{\alpha}^{(v)}{\bm{x}}|+\sqrt{2}|{\bm{D}}_{\alpha}^{(h)}{\bm{x}}|\right)^{2}\leq 4\left(({\bm{D}}_{\alpha}^{(v)}{\bm{x}})^{2}+({\bm{D}}_{\alpha}^{(h)}{\bm{x}})^{2}\right)$
Therefore, recall Eq. 36. It holds that
(38) $\begin{split}\left\|H_{ij}({\bm{x}})\right\|_{2}\leq~{}&\delta\sum_{\alpha\text{ in boundary}}\left(({\bm{D}}_{\alpha}^{(v)}{\bm{x}})^{2}+({\bm{D}}_{\alpha}^{(h)}{\bm{x}})^{2}+\varepsilon\right)^{-3/2}\\ &\cdot\left[4\left(({\bm{D}}_{\alpha}^{(v)}{\bm{x}})^{2}+({\bm{D}}_{\alpha}^{(h)}{\bm{x}})^{2}\right)+4\varepsilon\right]\\ \leq~{}&4\delta\sum_{\alpha\text{ in boundary}}\left(({\bm{D}}_{\alpha}^{(v)}{\bm{x}})^{2}+({\bm{D}}_{\alpha}^{(h)}{\bm{x}})^{2}+\varepsilon\right)^{-1/2}\leq 8m\delta\varepsilon^{-1/2}.\end{split}$
A similar result holds when $i,j$ are horizontal neighbors. Also notice $H_{jk}({\bm{x}})=0$ when $j\not\sim k$, where we denote $j\sim k$ if $j\neq k$ are neighboring blocks. Now substitute the above inequalities into Eq. 37. We have
(39) $\begin{split}&-\left(\nabla_{j}u({\bm{x}})\right)^{\rm T}\nabla_{j}\phi({\bm{x}}_{i})+\left(\nabla_{j}u({\bm{x}})\right)^{\rm T}\mathcal{L}(\nabla_{j}u)({\bm{x}})\\ \geq~{}&\lambda M_{jj}\|\nabla_{j}u({\bm{x}})\|_{2}^{2}-\sum_{k\neq j}(\lambda M_{jk}+8m\delta\varepsilon^{-1/2}{\bf{1}}_{j\sim k})\|\nabla_{j}u({\bm{x}})\|_{2}\|\nabla_{k}u({\bm{x}})\|_{2}.\end{split}$
Consider ${\bm{x}}$ where $\|\nabla_{j}u({\bm{x}})\|_{2}$ reaches its maximum, i.e. $\|\nabla_{j}u({\bm{x}})\|_{2}=\left\|\nabla_{j}u\right\|_{L^{\infty}}$. The first order optimality condition reads
$0=\nabla\left(\|\nabla_{j}u({\bm{x}})\|_{2}^{2}\right)=\nabla\nabla_{j}u({\bm{x}})\cdot\nabla_{j}u({\bm{x}}),$
@@
$0\geq\Delta\left(\|\nabla_{j}u({\bm{x}})\|_{2}^{2}\right)=2\|\nabla\nabla_{j}u({\bm{x}})\|_{\rm F}^{2}+2(\nabla_{j}u({\bm{x}}))^{\rm T}\Delta\nabla_{j}u({\bm{x}}).$
Thus, $(\nabla_{j}u({\bm{x}}))^{\rm T}\Delta\nabla_{j}u({\bm{x}})\leq 0$, since $\|\nabla\nabla_{j}u({\bm{x}})\|_{\rm F}^{2}\geq 0$. Under these conditions,
$(\nabla_{j}u({\bm{x}}))^{\rm T}\mathcal{L}(\nabla_{j}u({\bm{x}}))=(\nabla_{j}u({\bm{x}}))^{\rm T}\Delta\nabla_{j}u({\bm{x}})+\nabla\log\pi_{\varepsilon}({\bm{x}})\cdot\nabla\nabla_{j}u({\bm{x}})\cdot\nabla_{j}u({\bm{x}})\leq 0.$
Hence, at the maximum point Eq. 39 reads
$\lambda M_{jj}\|\nabla_{j}u\|_{L^{\infty}}^{2}-$ $\sum_{k\neq j}(\lambda M_{jk}+8m\delta\varepsilon^{-1/2}{\bf{1}}_{j\sim k})\left\|\nabla_{j}u\right\|_{L^{\infty}}\|\nabla_{k}u({\bm{x}})\|_{2}$
@@
Here, we use $\|\nabla_{j}\phi({\bm{x}}_{i})\|_{2}\leq\delta_{ij}$, since $\phi$ is $1$-Lipschitz and is only a function of ${\bm{x}}_{i}$. So if $\|\nabla_{j}u\|_{L^{\infty}}>0$, it holds that
@@
$\max_{\alpha}\mathbb{E}_{{\bm{x}}\sim\pi}\left\|{\bm{D}}_{\alpha}{\bm{x}}\right\|_{2}^{-1}\leq C_{\pi}.$
@@
${\bm{D}}_{\alpha}{\bm{x}}=({\bm{D}}_{\alpha}^{(v)}{\bm{x}},{\bm{D}}_{\alpha}^{(h)}{\bm{x}})=({\bm{x}}_{\alpha}-{\bm{x}}_{\alpha}^{(v)},{\bm{x}}_{\alpha}-{\bm{x}}_{\alpha}^{(h)}).$
Now introduce the change of variable ${\bm{z}}=T{\bm{x}}$ for some linear map determined via
${\bm{z}}_{\alpha}^{(v)}={\bm{x}}_{\alpha}^{(v)}-{\bm{x}}_{\alpha},\quad{\bm{z}}_{\alpha}^{(h)}={\bm{x}}_{\alpha}^{(h)}-{\bm{x}}_{\alpha},$
@@
and ${\bm{z}}_{\alpha-}={\bm{x}}_{\alpha-}\in\mathbb{R}^{d-2}$ for the other coordinates. Accordingly, $\pi({\bm{x}})$ is transformed into another distribution $\mu({\bm{z}})=\pi(T^{-1}{\bm{z}})$ (note $\det T=1$). Also note that $T^{-1}$ admits explicit form, i.e.
${\bm{x}}_{\alpha}^{(v)}={\bm{z}}_{\alpha}^{(v)}+{\bm{z}}_{\alpha},\quad{\bm{x}}_{\alpha}^{(h)}={\bm{z}}_{\alpha}^{(h)}+{\bm{z}}_{\alpha},\quad{\bm{x}}_{\alpha-}={\bm{z}}_{\alpha-}.$
Denote ${\bm{z}}_{\alpha+}=({\bm{z}}_{\alpha}^{(v)},{\bm{z}}_{\alpha}^{(h)})$ for convenience. Consider the factorization
$\mu({\bm{z}})=\mu({\bm{z}}_{\alpha+},{\bm{z}}_{\alpha-})=\mu({\bm{z}}_{\alpha+}|{\bm{z}}_{\alpha-})\mu({\bm{z}}_{\alpha-}),$
where $\mu({\bm{z}}_{\alpha-})$ denotes the marginal of $\mu$ on ${\bm{z}}_{\alpha-}$. Notice
@@
$\log\mu({\bm{z}}_{\alpha+}|{\bm{z}}_{\alpha-})=-\frac{\lambda}{2}\left\|{\bm{y}}-{\bm{A}}T^{-1}{\bm{z}}\right\|_{2}^{2}-\delta\sum_{\alpha=1}^{d}\left\|{\bm{D}}_{\alpha}(T^{-1}{\bm{z}})\right\|_{2}-\log\mu({\bm{z}}_{\alpha-})+\text{const}.$
@@
Fix ${\bm{z}}_{\alpha-}$ for the moment. Notice $\frac{\lambda}{2}\left\|{\bm{y}}-{\bm{A}}T^{-1}{\bm{z}}\right\|_{2}^{2}$ is $L$-smooth for some dimension-independent $L>0$, since only a dimension-independent number of coordinates of ${\bm{A}}T^{-1}{\bm{z}}$ depend on ${\bm{z}}_{\alpha-}$. Therefore, fix any ${\bm{z}}_{\alpha+}^{0}$, so that
$\frac{\lambda}{2}\left\|{\bm{y}}-{\bm{A}}T^{-1}({\bm{z}}_{\alpha+},{\bm{z}}_{\alpha-})\right\|_{2}^{2}-\frac{\lambda}{2}\left\|{\bm{y}}-{\bm{A}}T^{-1}({\bm{z}}_{\alpha+}^{0},{\bm{z}}_{\alpha-})\right\|_{2}^{2}-{\bm{v}}^{0}\cdot({\bm{z}}_{\alpha+}-{\bm{z}}_{\alpha+}^{0})$
$\leq\frac{L}{2}\left\|{\bm{z}}_{\alpha+}-{\bm{z}}_{\alpha+}^{0}\right\|_{2}^{2}$ ,
@@
where ${\bm{v}}^{0}$ is the gradient w.r.t. ${\bm{z}}_{\alpha+}$ of $\frac{\lambda}{2}\left\|{\bm{y}}-{\bm{A}}T^{-1}{\bm{z}}\right\|_{2}^{2}$ at ${\bm{z}}_{\alpha+}^{0}$. Notice also
$\delta\sum_{\alpha=1}^{d}\left\|{\bm{D}}_{\alpha}(T^{-1}({\bm{z}}_{\alpha+},{\bm{z}}_{\alpha-}))\right\|_{2}\leq\delta\sum_{\alpha=1}^{d}\left\|{\bm{D}}_{\alpha}(T^{-1}({\bm{z}}_{\alpha+}^{0},{\bm{z}}_{\alpha-}))\right\|_{2}+8\delta\left\|{\bm{z}}_{\alpha+}-{\bm{z}}_{\alpha+}^{0}\right\|_{2},$